### v3

# Why LLMs Are Bad at Synthetic Table Generation

Abstract.
Synthetic data generation is integral to ML pipelines, e.g., to augment training data, replace sensitive information, and even to power advanced platforms like DeepSeek. While LLMs fine-tuned for synthetic data generation are gaining traction, synthetic table generation—a critical data type in business and science—remains under-explored compared to text and image synthesis. This paper shows that LLMs, whether used as-is or after traditional fine-tuning, are inadequate for generating synthetic tables. Their autoregressive nature, combined with random order permutation during fine-tuning, hampers the modeling of functional dependencies and prevents capturing conditional mixtures of distributions essential for real-world constraints. We demonstrate that making LLMs permutation-aware can mitigate these issues. Our code and data are anonymously hosted [^1] https://github.com/ShengzheXu/Permutation-aided-Fine-tuning.
copyright: acmlicensedjournalyear: 2025
1. Introduction
Large language models (LLMs) have found applicability in a rich diversity of domains, well beyond their original roots (Achiam et al., 2023; Brown et al., 2020; Miller and Blair, 2009; OpenAI, 2022; Touvron et al., 2023; Raffel et al., 2020; Romera-Paredes et al., 2023). As so-called foundation models (Liang et al., 2022) they have been shown to be re-targetable to a variety of downstream tasks. Our focus here is to view LLMs as raw synthetic tabular data generators rather than as supporting an analysis or discovery task. Arguably, LLMs are adept at synthetic generation of text, images, videos, code, documentation, and many other modalities. The use of LLMs to generate tabular data is quite understudied.
Such synthetic tabular data generation is integral to ML pipelines, e.g., too augment training data, replace sensitive information, and even to power advanced platforms like DeepSeek (Liu et al., 2024; Guo et al., 2025). However, the unique characteristics of tabular data manifest as challenges in an LLM-based generative context. The most popular incarnations of language models are auto-regressive models, e.g., LLama (Touvron et al., 2023), GPT-x (OpenAI, 2022), DeepSeek (Liu et al., 2024; Guo et al., 2025), wherein each word or token is generated conditional on past tokens in a sequential manner using attention models. In a synthetic data context, each ‘sentence’ typically represents a row of tabular data, and each ‘word’ corresponds to an attribute in that row. The previous state-of-the-art models (GReaT (Borisov et al., 2022)), has advocated the use of random feature orders but as we show in Table 1, when fine-tuning is done with random feature orders, key relationships are often not captured or, worse, violated. In particular, with tabular data, there are numerous functional dependencies at play and as a result, generating tokens in random orders is bound to cause violations. There is thus an ‘impedance mismatch’ between autogressive LLMs and synthetic data generation.
The main contributions of this paper are:
  • We highlight an important deficiency with using LLMs for synthetic tabular data generation and explore the performance of many state-of-the-art generation models in the context of composite and multi-category tabular schema.
  • We inject knowledge of pre-existing functional relationships among columns into the autoregressive generation process, so that the generated synthetic data respects more real constraints. In particular, we present a taxonomy of functional dependencies (FDs) whose discovery and organization into a column dependence graph supports their incorporation into the LLM fine-tuning process via a permutation function, leading to our approach dubbed Permutation-aided Fine-tuning.(PAFT ).
  • We evaluate the performance of PAFT on a range of datasets featuring a diverse mix of attribute types, functional dependencies, and complex relationships. Our results demonstrate that PAFT is the state-of-the-art in reproducing underlying relationships in generated synthetic data.
  • Finally, we demonstrate through rigorous experiments that relying just on standard univariate distribution, bivariate correlation, and even the evaluation of downstream machine learning models (which primarily focuses on predicting a single column in a dataset) is grossly insufficient for assessing the quality of synthetic data and propose systematic remedies like measuring violation rates of known domain rules.
2. Related Work
Table 1. Comparison of multiple synthetic data generation approaches. The second row showcases the state of West Virginia (WV), which is a subset of the whole data. In the figures, the solid line represents the official border of West Virginia and the Blue and Orange colors indicate the legal and illegal samples in the synthetic data respectively.
Whole Data
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Sub-Category
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Real CTGAN CopulaGAN TabSyn (Diffusion) GReaT (LLM-GPT) PAFT (LLM-GPT)
Tabular data synthesis and representation learning for tables have been extensively studied (Du et al., 2021; Wang et al., 2021; Zhang et al., 2023b; Margeloiu et al., [n. d.]; Fang et al., 2024; Patki et al., 2016; Du and Li, 2024; McKenna et al., 2019; Park et al., 2018; Wang et al., 2020). One key theme has been the discovery of functional dependencies (FDs) which has been studied by the data mining community from both theoretical and application points of view (Zhang et al., 2020; Chen et al., 2019; Muralidhar et al., 2018; Mandros et al., 2020; Pennerath et al., 2020; Zheng et al., 2023). We carve out related work into two sections, for our purposes: pre-LLM (or non-LLM) approaches for synthetic table generation which continue to hold the mainstay, and LLM approaches.
Lei et al. (Xu et al., 2019) proposed CTGAN where rows are independent of each other; a conditional GAN architecture ensures that the dependency between columns is learned. Tabsyn (Zhang et al., 2023b) showcased remarkable advancements in joint-distribution learning via a VAE plus diffusion approach, surpassing previous models of similar lineage, in terms of distributional correlation measures and machine learning efficiency. DoppelGanger (Lin et al., 2020) uses a combination of an RNN and a GAN to incorporate temporal dependencies across rows but this method has been tested in traditional, low-volume settings such as Wikipedia daily visit counts. For high-volume applications, STAN (Xu et al., 2021) utilizes a combination of a CNN and Gaussian mixture neural networks to generate synthetic network traffic data. GraphDF  (Chen et al., 2023) conducts multi-dimensional time series forecasting. GOGGLE  (Liu et al., 2022) employs a generative modeling method for tabular data by learning relational structures.
The use of language models (LLMs) for tabular data generation is still underexplored. Most modern LLMs are based on the transformer architecture  (Vaswani et al., 2017) with parameters ranging from few millions to billions  (Hoffmann et al., 2022), and researchers have developed creative ways to harness LLMs in traditional machine learning and data contexts. LIFT (Dinh et al., 2022) initially transforms a table row into a sentence, such as ‘An Iris plant with sepal length 5.1cm, sepal width 3.5cm’, and employs an LLM as a learning model for table classification, regression, and generation tasks. GReaT (Borisov et al., 2022) utilizes a GPT-2 model that has been fine-tuned using a specific corpus for synthetic data generation. They also show that even small-scale models such as Distill-GPT (Radford et al., 2019) have the potential for synthetic data generation (Borisov et al., 2022). These models are specially viable for tabular generation given the lower compute costs of aligning smaller models to large and varied tabular datasets. A general benefit of utilizing LLMs is the promise of eliminating customized preprocessing pipelines.
A theme that will be pertinent to our work is the idea of feature ordering (FO) for tabular data generation which has been investigated from multiple angles (Sharma et al., 2024; Distiawan et al., 2018; Ribeiro et al., 2019; Zhu et al., 2022). There are also several approaches (e.g., (Kanter and Veeramachaneni, 2015; Cvitkovic, 2020)) that synthesize, discover, or aggregate features from relational databases, leveraging order information when possible, for use in machine learning pipelines. It worth noting that even in the LLM community, the task of context sorting for LLM prompting is not trivial and has gained significant attention lately (Chen et al., 2024).
3. Challenges to Synthetic Table Generation in the Current LLM Paradigm
Figure 1. Overview of the proposed Permutation-Aided Fine-tuning (PAFT ) approach.
Admittedly, the sure-fire way to check if a dataset has been faithfully modeled is to see if the joint distribution of its features is captured accurately. Most current tests of synthetic data generation quality focus on fidelity to single-column distributions, or to multi-column distributions. For multiple variables, measures such as machine learning efficiency (MLE) (Borisov et al., 2022; Xu et al., 2019; Zhang et al., 2023b; Dinh et al., 2022) are frequently used to construct classifiers or regressors.
A key lesson from probabilistic graphical model research (Koller and Friedman, 2009) is that factorizing joint distributions into products of conditionals (e.g., Bayesian networks) dramatically helps reduce the number of parameters necessary to capture the underlying data characteristics. A similar lesson from database research (Garcia-Molina et al., 2008) is that modeling functional dependencies (FDs) in data helps reduce redundancy in modeling and storage. These lessons, i.e., that order matters, continue to apply in the LLM era but are not internalized in our prompt ordering, fine-tuning, or evaluation methodologies. Other researchers have noted the importance of ordering in LLMs (Chen et al., 2024; Prystawski et al., 2024) but this lesson has not been leveraged to improve synthetic table generation by LLMs. In the absence of leveraging good feature orders, existing approaches either focus on ‘one order’, ‘no order’, or ‘all orders’. All of these approaches severely limit the quality of generated synthetic data.
As an example, Table 1 presents a case study on using models to generate synthetic data of locations in various states of the USA. The data is of the form (state, latitude, longitude) where the attributes adhere to the FD: $\{\text{latitude},\text{longitude}\}\rightarrow\text{state}$. Existing methods can generate satisfactory univariate distributions (as shown in the border of the top row plots) but fail to capture the joint distribution (center of the top row plots) and conditional distributions across subcategories (bottom row plots).
In summary, feature ordering can be both a nuisance and a gift. It is a nuisance because it demands additional constraints to be modeled. It can be a gift because it suggests ways to sequentially generate data even by autoregressive LLMs. Our proposed approach (PAFT ) aims to achieve an optimal permutation order for fine-tuning LLMs.
Figure 1 shows an overview of the proposed permutation aided fine-tuning approach (PAFT ). A typical workflow is 1) Textual Encoding (Section 4.1) 2) Functional Dependency (FD) discovery (Section 4.2) 3) FD Distillation ( Section 4.2) and 4) Feature Order Permutation Optimization ( Section 4.3). Our fine-tuning and sampling strategy is explained in Section 4.4.
4. PAFT : Permutation-Aided Fine-Tuning
Problem Setup. Let $\mathcal{D}$ represent a table with $n$ rows (i.e., records) and $m$ columns (i.e., attributes a.k.a schema). Let each record be represented by vector $\mathbf{x}_{i}$ and further let $x_{ij}$ represent the element value of the $j^{th}$ attribute of record $\mathbf{x}_{i}$. Hence each row $\mathbf{x}_{i}\in\mathcal{D}$ represents an individual record and each column $\mathbf{x}_{(:,j)}\sim\mathcal{X}_{j}$ can be considered sampled from a random variable $\mathcal{X}_{j}$ that governs the distribution of attribute $j$. Finally, let $i\in[1..n]$ and $j\in[1..m]$. Realistically, tabular data $\mathcal{D}$ is frequently a mixture of categorical and continuous attributes, hence each $\mathcal{X}_{j}$ can be a categorical or continuous random variable. If $\mathcal{A}=\{\mathcal{X}_{1},\mathcal{X}_{2},\dots,\mathcal{X}_{m}\}$ represents the collection of random variables, then the table generation process aims to sample from a joint distribution $\mathbb{P}(\mathcal{A})=\mathbb{P}(\mathcal{X}_{1},\mathcal{X}_{2},\dots,\mathcal{X}_{m})$. This joint distribution is usually a complex, high-dimensional distribution and, most importantly, unknown. The goal of learning an effective tabular data generator $p_{\theta}(\cdot)$ is to enable $p_{\theta}(\cdot)$ to learn a faithful approximation $\mathbb{P}(\mathcal{A}|\mathcal{D})$ of the data generation process distribution $\mathbb{P}(\mathcal{A})$ using the data sample $\mathcal{D}$ such that $\mathbb{P}(\mathcal{A}|\mathcal{D})\approx\mathbb{P}(\mathcal{A})$. Once such an effective model $p_{\theta}(\mathcal{D})$ is trained, it can be employed to generate large volumes of seemingly realistic synthetic data $\mathcal{\hat{D}}\sim\mathbb{P}(\mathcal{A}|\mathcal{D})$.
4.1. Tabular Data Generation with LLMs
While training $p_{\theta}(\cdot)$, it is usually assumed that all records $\mathbf{x}_{i}\in\mathcal{D}$ are independent. Generating new data samples $\hat{\mathbf{x}}_{i}\in\hat{\mathcal{D}}$ can be done in various ways (e.g., see (Xu and Veeramachaneni, 2018; Xu et al., 2019; Zhao et al., 2021)) which aim to directly estimate the joint distribution $\mathbb{P}(\mathcal{A})$ or, as is done here in PAFT , where $\mathbb{P}(\mathcal{A})$ is estimated by an autoregressive LLM based generative process, as a product of multiple conditional densities governed by the input ordering.
Autoregressive LLM models are pre-trained to maximize the likelihood of target token $x_{ij}\in\mathcal{D}$, conditioned upon the autoregressive context $\mathbf{x}_{(i,1:j-1)}\in\mathcal{D}$ where $\mathcal{D}$ is the training corpus comprising a large amount of textual data (in the pre-training context). Eq. 1 defines the general training criterion of LLM training using the self-supervised next-token prediction task with ‘w’ denoting the context length.
(1) $\mathcal{L}(\theta;\mathcal{D})=-\sum\limits_{\mathbf{x}_{i}\in\mathcal{D}}\sum\limits_{j=1}^{w}\mathrm{log}\,\mathbb{P}(x_{i}|\mathbf{x}_{(i,1:j-1)}).$
The generation of a single instance (i.e., database record) $\mathbf{x}_{i}\in\mathcal{D}$ is given by Eq. 2:
(2) $\mathbb{P}(\textbf{x}_{i})=\mathbb{P}(x_{i,1},...,x_{i,m})\simeq\prod_{j=1}^{m}\mathbb{P}(x_{i,j}|x_{i,1},..,x_{i,j-1}).$
Specifically, each database record is generated as a product of conditional distributions.
Input Encoding. To support the processing of our records $\mathbf{x}_{i}\in\mathcal{D}$ by a pre-trained LLM, we adopt the following encoding:
(3) $\begin{split}t_{i,j}=[c_{j},\textrm{`}is\textrm{'},x_{i,j},\textrm{`},\textrm{'}],&i\in\{1,..,n\},j\in\{1,..,m\},\\ \mathbf{t}_{i}=[t_{i,1},t_{i,2},..,t_{i,m}],&i\in\{1,..,n\}.\end{split}$
In Eq. 3, $c_{j}$ represents the attribute name of the $j^{th}$ database column while $x_{i,j}$ represents the actual value of the $j^{th}$ column for the $i^{th}$ record. Further, we can assume we have a mechanism to obtain a feature order permutation $\mathbf{k}$ to govern the order of the attributes in $\mathbf{t}_{i}$, such that $\mathbf{t}_{i}(\mathbf{k})=[t_{i,k_{1}},t_{i,k_{2}},..,t_{i,k_{m}}]$ (where $i\in\{1,..,n\},k_{j}\in\{1,..,m\}$.), represents the same record but with the attribute order governed by the permutation $\mathbf{k}$, This definition admits the random feature order as a special case in which $\mathbf{k}$ is a random permutation.
Since we consider autoregressive LLM-based generative models, employing the chain rule to sequentially produce each column of a table record $\mathbf{t}_{i}$, we can view each generation step as approximating the joint distribution of the table columns as a product of conditional distributions (i.e., $\mathbb{P}(t_{i,1},...,t_{i,m})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,j}|t_{i,1},..,t_{i,{j-1}})$). However, as the number of columns increases and the relationships between columns get more conditional, the likelihood of encountering training and sampling bias due to class imbalance also rises (Xu et al., 2019). To minimize such adverse effects, we can consider injecting knowledge of the pre-existing functional relationships among columns, to govern the autoregressive generation process. To infer such functional relationships, we leverage a learned dependency graph derived from functional dependency (FD) relations which enables us to effectively determine the appropriate training and sampling sequence. This, in turn, allows us to alleviate potential biases during training by establishing a generation curriculum leading to improved estimation accuracy of the joint distribution $\mathbb{P}(\textbf{t}_{i})$ in auto-regressive prediction $\mathbb{P}(t_{i,k_{1}},...,t_{i,{k_{m}}})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,k_{j}}|t_{i,k_{1}},..,t_{i,k_{j-1}})$, where the ordering $t_{i,k_{1}}\dots,t_{i,k_{m}}$ is obtained by a feature order permutation function $\mathbf{k}=\phi(\mathcal{D},\mathcal{S})$. We detail the requisite background and design of $\phi(\mathcal{D},\mathcal{S})$ in sections 4.2 and  4.3.
4.2. Discovery and Distillation of Functional Dependencies (FD)
A functional dependency (FD) is a relationship $R$ in schema $S$ that exists when a subset of attributes $A\subset S$ uniquely determines another subset $B\subset S$ of attributes. We succinctly represent an FD as $R:A\rightarrow B$ which specifies that $B$ is functionally dependent on $A$.
Definition 0 (Schema-Level FD).
With $A$, $B$ being two disjoint subsets of the schema (columns) of table $\mathcal{D}$, a schema-level FD $F$ associated with $\mathcal{D}$ has the form: $F:A\rightarrow B$.
We leverage FD discovery techniques to govern the order of the autoregressive data generation process in PAFT . A large body of research from the database literature on FD discovery (Papenbrock et al., 2015; Zhang et al., 2020; Papenbrock and Naumann, 2016) can be leveraged in PAFT , including methods that account for noisy FDs (Zhang et al., 2020). In this work, we focus on leveraging schema-level FDs discovered using a state-of-the-art FD discovery algorithm (Papenbrock and Naumann, 2016) to govern the autoregressive data generation process.
FD Distillation. The result of traditional FD discovery, yields complex (i.e., multi-attribute) functional dependencies between columns which are ambiguous to resolve in an autoregressive generation setting. Hence we undertake an intermediate FD distillation step to simplify multi-attribute functional dependencies into multiple single attribute FDs as detailed below.
(a) There are two types of column dependency edges for three types of functional dependencies (FDs), which are distinguished by the left-hand side (LHS) and right-hand side (RHS) in the FD.
(b) DAG for column functional dependency derived by expanding SCC super nodes and retrieving a fully flattened, ordered structure.
Figure 2. FD-Distillation and Dependency Graph Sorting for automatically extracting order permutations from tables.
We first construct a dependency graph model $G(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ represents the set of vertices with each $v\in\mathcal{V}$ representing an attribute in $\mathcal{S}$ and $e_{ij}\in\mathcal{E}$ representing an edge relation from attribute $v_{i}$ to attribute $v_{j}$. Further, we consider two types of edges in $\mathcal{E}$, specifically each $e_{ij}$ may be a type-1 edge or a type-2 edge (defined next). The two edge types (i.e., type-1, type-2) in $\mathcal{E}$ are derived from three classes of FDs, as shown in Fig. 2(a). Subsequently, we proceed to examine each individual case: 1) Single attribute left-hand side (LHS) and single attribute right-hand side (RHS) 2) Multi-attribute LHS and single-attribute RHS. 3) single or multi-attribute LHS and multi-attribute RHS. We shall use the example table in Fig.  1 to define each FD case.
[Type-1 Edge]. Let us consider an example of FD Case 1, wherein the column State functionally determines column Bird. In such a case, we enforce that the value for the attribute State be generated prior to the value for the attribute Bird. Accordingly, a forward directed edge from State to Bird is created in the column dependency graph $\mathcal{G}$. We term such forward directed edges as type-1 edges in $\mathcal{G}$. [Type-2 Edge]. The other type of edge in $\mathcal{G}$, arises when we encounter an FD with a multi-attribute LHS and a single attribute RHS (i.e., FD Case 2). As per FD Case 2, the values of multiple columns in the LHS would collectively decide the value of the column on the RHS. As an example in Fig. 2(a), the tuple of columns Latitude, Longitude functionally determines State. For such FDs, two backward edges are added in the column dependency graph $\mathcal{G}$, connecting State to both Latitude and Longitude. We term such backward directed edges as type-2 edges in $\mathcal{G}$.
FD Case 3 relationships are ones where the RHS has multiple attributes and the LHS could have single or multiple-attributes. Such relationships do not directly result in an edge in our dependency graph $\mathcal{G}$. Instead, as classically done in FD literature (Papenbrock and Naumann, 2016), we subject such FD relationships to an intermediate decomposition step. Specifically, the multi-attribute RHS of FD Case 3 relationships is decomposed into multiple single-attribute RHS dependencies each comprising the original LHS. Further, in each of these new decomposed relationships, if the LHS is single-attribute, it is treated as a FD Case 1 relationship (i.e., a directed edge from $LHS$ to $RHS$ is added to $\mathcal{G}$), else it is handled as an FD Case 2 relationship, wherein for each attribute in the multi-attribute LHS, a backward dependency edge from the single-attribute RHS is added to the dependency graph $\mathcal{G}$. Therefore, for every column in the right-hand side (RHS), we employ either type-1 or type-2 edge construction, depending on the value of its left-hand side (LHS).
4.3. Putting It All Together
Until this point, our construction of the graph $\mathcal{G}$ has only been limited to considering pair-wise relationships between columns in $\mathcal{D}$. Graph properties like functional dependency transitivity, require us to obtain a total ordering on the nodes $v\in\mathcal{V}$ of $\mathcal{G}(\mathcal{V},\mathcal{E})$ that is deterministic in nature for effective auto-regressive LLM training. This implies that in order to obtain a feature order permutation ($\mathbf{k}$) from the derived functional dependency relationships (Sec. 4.2), a computation must be performed on the entire dependency graph $G(\mathcal{V},\mathcal{E})$.
We define this task of obtaining a total feature order $\mathbf{k}$ from $\mathcal{G}(\mathcal{V},\mathcal{E})$ as an optimization step which seeks to produce $\mathbf{k}$ while minimizing the number of violated relationships in $\mathcal{G}(\mathcal{V},\mathcal{E})$.
Consider the trivial case of having an empty FD graph $\mathcal{G}$ i.e., $|\mathcal{E}|=\emptyset$. If the columns of a table are not functionally dependent on each other, then the order of generation is not important and $\mathbf{k}$ can be some arbitrary permutation of the columns in $\mathcal{S}$. For all other cases, our total feature ordering algorithm operates in three phases, as shown in Fig. 2(b). [Phase 1: Condensation]. It is apparent that if a graph is not a directed acyclic graph (DAG), there is no optimal solution to the total feature ordering problem. In other words, there must be FDs that cannot be satisfied in the resulting total order permutation $\mathbf{k}$. In such cases, we compute the strongly connected components (SCC), and condense them into super nodes, thus transforming the original graph into a DAG. [Phase 2: Ordering]. An application of a topological sort onto the DAG from Phase 1 will result in a total feature ordering with all SCCs in $\mathcal{G}$ compressed into super nodes. [Phase 3: Expansion of SCC]. Once the topological sort is conducted in Phase 2, we finally expand the SCC super nodes (via. arbitrary ordering) such that although the intra-SCC ordering of the nodes within the SCC is arbitrary, their ordering relative to non-SCC nodes is maintained.
4.4. Synthetic Data Generation using PAFT
After the optimized feature order permutation is obtained, we fine-tune the LLM with the textually encoded table record $\mathbf{t}_{i}$ such that the auto-regressive generation process is governed by the optimal feature order permutation $\mathbf{k}=\phi(\mathcal{D},S)$. Specifically, we generate the table governed by order $\mathbf{k}$ as defined in Eq. 4.
(4) $\mathbb{P}(\textbf{t}_{i})=\mathbb{P}(t_{i,k_{1}},...,t_{i,{k_{j}}})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,k_{j}}|t_{i,k_{1}},..,t_{i,k_{j-1}}).$
We employ the Low-Rank Adaptation (LoRA) fine-tuning strategy (Hu et al., 2021). To generate synthetic rows, we first sample the initial token $p(t_{i,k_{1}})$ from the marginal distribution of variable $k_{1}$ in actual training data, and then use Eq. 4 to sequentially sample subsequent tokens $p(t_{i,k_{j}})$, where $j\in{2,...,m}$.
5. Experimental Evaluation
We conduct an exhaustive empirical evaluation of PAFT to assess its ability to reproduce realistic data distributions, superiority over other competing approaches, and most importantly substitutability of data generated by PAFT in the context of a larger ML pipeline. More specifically, the questions we seek to answer are:
  1. (1)
    Does PAFT-generated synthetic data accurately capture conditional distributions within categories? (Sec 5.1)
  2. (2)
    Does PAFT generate data respect the consistency of intrinsic data characteristics? (Sec 5.2)
  3. (3)
    Does the synthetic data generated by PAFT pass the sniff test? (Sec 5.3)
  4. (4)
    Can data generated by PAFT replace real data in downstream ML model training? (Sec 5.4)
  5. (5)
    Do the data sets generated by PAFT adhere to real distributions and possess mode diversity? (Sec 5.5)
  6. (6)
    Do newer generations of LLMs obviate the need for PAFT ? (Sec 5.6)
Datasets. We evaluate the effectiveness of PAFT through experiments on six real datasets commonly used in synthetic table generation studies such as GReaT (Borisov et al., 2022) CTGAN (Xu et al., 2019). These are Beijing (Chen, 2017), US-locations (Evans-Bye, 2015), California Housing (Nugent, 2017), Adult Income (Barry Becker, 1996), Seattle (samuel Cortinhas, 2022), and Travel (Tejashvi, 2021). Separately, we also generate a set of four simulated datasets for class-mixture distributions.
Baselines. For benchmarking, we organize baselines that utilize current deep learning approaches for synthetic data generation (CTGAN (Xu et al., 2019), CopulaGAN (Xu et al., 2019), TabSyn (Zhang et al., 2023b)) and the most advanced synthetic table generator with LLM fine-tuning GReaT (Borisov et al., 2022). To guarantee an equitable comparison, we employ the Distill-GReaT model for both LLM techniques in all tests.
Reproducibility. Each baseline (CTGAN, CopulaGAN, TabSyn, GReaT) adheres to the recommended hyperparameters and utilizes officially released API tools: Synthetic Data Vault (Patki et al., 2016) and GReaT (Borisov et al., 2022). For a fair comparison of GReaT and PAFT, the LoRA fine-tuning parameters are set the same as: Lora attention dimension $r=16$, alpha parameter for Lora scaling $lora\_alpha=32$, the names of the modules to apply the adapter to $target\_modules=c\_attn$, the dropout probability for Lora layers $lora\_dropout=0.05$, $bias=none$.
Parameters for MLE and Discriminator Models. We utilize neural network, linear/logistic regression, and random forest models from the Scikit-Learn package for the ML efficiency and discriminator experiments. Every result is evaluated through the 5-fold cross-validation process.
Low-order statistics (Zhang et al., 2023b) of column-wise data distribution and pair-column correlation are calculated with the SDV library  [^2] https://docs.sdv.dev/sdmetrics/reports/quality-report.
5.1. RQ1: Does PAFT-generated synthetic data accurately capture conditional distributions within categories?
Figure 3. For a composite dataset, US-locations, this comparison examines state-specific violation rates across different synthetic data generation approaches. The error bars represent standard deviation. The states on the x-axis are ordered by decreasing violation rates. PAFT significantly reduces state-specific violations in the composite dataset.
Table 2. Violation rates across different categories in the same dataset highlight the effectiveness of PAFT in addressing conditional distribution challenges in mixed-category datasets. Notably, even though baseline models may perform well in the MLE task or distribution evaluation, they often fail in practical boundary checks, such as functional dependencies (FDs). We can see that states with the lowest violation rates are the easiest to model, i.e. rectangular.
States with the Highest Violation Rates ($\downarrow$) (Sorted by LLM baseline: GReaT)
Category (State) MO AK KY FL WV
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Real Data 0% 0% 0% 0% 0%
CTGAN 98.0% 99.3% 97.4% 99.9% 100%
CopulaGAN 99.3% 84.6% 99.3% 97.8% 100%
TabSyn 10.5% 17.2% 22.9% 25.1% 28.1%
GReaT 17.1% 18.4% 20.6% 21.9% 22.3%
PAFT 1.9% 5.1% 3.4% 3.4% 3.4%
States with the Lowest Violation Rates ($\downarrow$) (Sorted by LLM baseline: GReaT)
Category (State) CO KS NM SD UT
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Real Data 0% 0% 0% 0% 0%
CTGAN 96.5% 97.0% 99.5% 97.5% 99.3%
CopulaGAN 98.3% 98.8% 98.4% 98.3% 98.4%
TabSyn 11.3% 18.4% 6.2% 17.3% 16.6%
GReaT 0.5% 0.9% 1.2% 1.5% 2.1%
PAFT 0.0% 0.3% 0.9% 0.9% 0.0%
Capturing and generating the diversity in a multi-class setting (composite dataset) has been shown to be a challenging practical problem (Margeloiu et al., [n. d.]). In addition to Table 1, Fig. 3 and Table 2 further illustrate the widespread nature of this issue within the same composite dataset, where the GAN-generated results exhibit an almost 100% violation rate. The term ‘composite dataset’ refers to a dataset in which the distributions across different subcategories show significant variances. The LLM model GReaT, which employs random order permutation, shows a significant improvement in maintaining conditional distributions. Nevertheless, its performance remains inconsistent due to unevenness within categories, resulting in violation rates ranging from 0.5% to 22.3%. In contrast, PAFT consistently controls deviations from the real facts of conditional distributions, maintaining them within a range of 0% to 5%. This reliability holds even in challenging subcategory cases where baseline methods underperform.
5.2. RQ2: Does PAFT generate data respecting the consistency of intrinsic data characteristics?
Table 3. Datasets have intrinsic characteristics like functional dependencies, range restrictions, and other domain knowledge. Results are averaged over five random runs.
Intrinsic Fact (Dataset) Fact Violation Rate ($\downarrow$)
CTGAN Cop.GAN TabSyn GReaT PAFT
[…]
Lat-long $\rightarrow$ CA (California) 47.6% 99.9% 8.8% 5.4% 1.3%
[…]
$\rightarrow$ education-num (Adult) 83.9% 19.1% 1.4% 1.2% 0.5%
[…]
In addition to the dissimilar sub-category challenge, we also investigate whether unsatisfactory conditional distributions exist in general across various datasets, and whether the PAFT method can address these issues. In line with this, we conducted rule checks that were derived from real-world scenarios and subsequently evaluated the generated data from all models. Table 3 displays the violation rate in the generated data. From the table, we observe that PAFT adheres to data’s characteristics more faithfully (i.e., significantly fewer rule violations) than baseline methods, by learning together with functional dependencies.
5.3. RQ3: Does the synthetic data generated by PAFT pass the sniff test?
Similarly to the analysis conducted in recent work (Borisov et al., 2022) (GReaT), we employ the random forest (RF) algorithm to train discriminators to distinguish real data (labeled True) and synthetically generated data (labeled False). Subsequently, we test performance on an unseen set (consisting of 50% synthetically generated data and 50% real data). In this experiment, scores represent the percentage of correctly classified entities. In this case, an ideal accuracy score would be close to 50%, which means the discrimniator fails to distinguish between real and synthesized data. The scores are shown in Table 4 and indicate that the data generated by PAFT are most indistinguishable from real data, even by powerful discriminative models.
Table 4. Privacy Discriminator Performance. The scores stand for the accuracy for detecting real or fake data, where the ML models are trained using 50% real data and 50% random data. An ideal accuracy score is 50, indicating the model cannot distinguish between real and synthesized data. The best results are marked in bold, the second-best results are underlined. Results are averaged over five random runs.
Data Sniff Test - ML Discriminator Accuracy (Values closest to 50% are best.)
Method CTGAN Co.GAN TabSyn GReaT PAFT
Beijing 99.16% 98.69% 50.97% 51.1% 50.09%
US-locations 99.94% 97.74% 51.97% 50.47% 50.01%
California 98.35% 86.64% 50.64% 53.74% 49.89%
Adult 94.43% 59.82% 51.64% 51.12% 48.75%
Seattle 87.61% 85.7% 50.12% 68.27% 47.21%
Travel 77.96% 74.14% 50.66% 62.49% 48.18%
Table 5. MLE performance. MLE performance: For datasets with regression tasks (marked *), performance is evaluated using MAPE (where lower scores are better). For datasets with classification tasks, accuracy is used (where higher scores are better). The best results are marked in bold and the second-best results are underlined. Results are averaged over five random runs
. Regression MAPE ($\downarrow$ Over Different Methods Dataset(*) Orig. CTGAN Co.GAN TabSyn GReaT PAFT Beijing RF 0.41% 2.49% 2.15% 0.7% 0.57% 0.52% LR 1.37% 2.23% 1.55% 1.25% 0.97% 1.34% NN 0.99% 2.44% 2.83% 1.01% 1.16% 0.95% Calif. RF 0.18% 0.65% 0.39% 0.22% 0.25% 0.20% LR 0.30% 0.54% 0.5% 0.30% 0.29% 0.31% NN 0.34% 0.53% 0.47% 0.29% 0.3% 0.27% Seattle RF 0.33% 0.76% 0.38% 0.30% 0.35% 0.28% LR 0.29% 0.74% 0.32% 0.23% 0.33% 0.29% NN 0.28% 0.71% 0.38% 0.28% 0.33% 0.27% Classif. Accuracy ($\uparrow$ Over Different Methods Dataset Orig. CTGAN Co.GAN TabSyn GReaT PAFT US-loc. RF 99.95% 7.17% 45.33% 99.99% 99.84% 99.91% LR 46.1% 5.11% 31.08% 43.69% 45.65% 49.41% NN 99.85% 7.56% 53.34% 99.64% 98.94% 99.44% Adult RF 84.97% 71.15% 81.33% 83.69% 83.89% 83.06% LR 78.53% 75.68% 78.18% 78.38% 76.1% 77.24% NN 76.9% 75.69% 76.6% 78.36% 78.23% 79.16% Travel RF 88.95% 56.35% 67.18% 84.09% 79.78% 85.19% LR 82.87% 70.17% 79.56% 83.31% 78.34% 82.76% NN 81.77% 71.05% 79.56% 81.88% 80.77% 83.20%
5.4. RQ4: Can data generated by PAFT replace real data in downstream ML model training?
We next assess the effectiveness of the generated (synthetic) data by comparing the performance of discriminative models trained on synthetic data versus real data for their target tasks. Models tested include random forests (RF), linear regression (LR), and multi-layer perceptron (NN). As shown in Table 5, PAFT  is best or second best in over 80% of (dataset, method) combinations.
5.5. RQ5: Do the data sets generated by PAFT adhere to real distributions and possess mode diversity?
We also evaluate how closely the density and diversity of the true data distribution are matched by PAFT generated data using correlation metrics and density-based distance metrics. Specifically, we employ the Kolmogorov-Smirnov Test (KST) to evaluate the density estimate of numerical columns, and the Total Variation Distance (TVD) for categorical columns. When calculating the correlation between columns, we employ Pearson correlation for numerical columns and contingency similarity for categorical columns. The results for both density estimate similarity and correlation based analysis are detailed in Table 6. As shown, PAFT synthetic data closely matches the real data in terms of univariate distribution and bivariate correlation, outperforming the baseline.
Table 6. Low-order statistics (Zhang et al., 2023b) of column-wise data density and pair-wise column correlation2. Scores range from 0 to 1. Higher values indicate more accurate estimation. PAFT outperforms the best generative baseline model in most case. The best results are marked in bold, the second-best results are underlined. Results are averaged over five random runs.
Single-Column Shape Score ($\uparrow$)
Dataset CTGAN Co.GAN TabSyn GReaT PAFT
Adult 0.81 0.92 0.98 0.88 0.90
Beijing 0.89 0.79 0.98 0.93 0.97
California 0.87 0.77 0.98 0.89 0.83
US-locations 0.83 0.82 0.96 0.93 0.97
Seattle 0.83 0.73 0.93 0.90 0.93
Travel 0.84 0.90 0.93 0.93 0.93
Two-Column Pair Trends score ($\uparrow$)
Dataset CTGAN Co.GAN TabSyn GReaT PAFT
Adult 0.81 0.86 0.93 0.80 0.78
Beijing 0.92 0.94 0.99 0.95 0.98
California 0.84 0.87 0.97 0.87 0.91
US-locations 0.50 0.55 0.93 0.89 0.94
Seattle 0.74 0.72 0.80 0.76 0.81
Travel 0.77 0.80 0.87 0.85 0.82
Visual examples are depicted in Figure 4. PAFT has the ability to generate a wide range of diversity, encompassing both continuous and discrete variables, which closely resembles real data.
Figure 4. Column distributions visualization for each dataset generated by CTGAN, CopulaGAN, GReaT, and PAFT . The top row displays examples of numerical columns, while the bottom row presents examples of categorical columns. Overall, PAFT (Blue) has the closest distribution to real data (Red) compared to other synthesis methods. PAFT also showcase the ability to generate a wide range of diversity.
5.6. RQ6: Do newer generations of LLMs obviate the need for PAFT ?
The choice of foundation models for synthetic table generation involves two key considerations: in-context learning and fine-tuning, especially with large-parameter models such as GPT-4 and LLaMA. A recent survey (Fang et al., 2024) indicates that the in-context learning approach of GPT-4 is well-suited for augmenting tabular data in low-data regimes, as highlighted by CLLM (Seedat et al., 2024). However, GPT-4 suffers from limitations such as the disappearance of column-wise tail distributions and a low success rate in accurately extracting output cell values (Fang et al., 2024).
In contrast, the state-of-the-art GReaT (Borisov et al., 2022) and other synthetic data generation approaches (Solatorio and Dupriez, 2023; Zhang et al., 2023a; Zhao et al., 2023) typically employ smaller models like GPT-2 or DistilGPT2, which effectively address attribute encoding while reducing feature engineering efforts. These smaller models already outperform traditional GAN and traditional statistics-based approaches. For example, DistilGPT2 can be trained using LoRA on GPUs as modest as the Tesla P100.
Importantly, fine-tuning larger models such as GPT-4 entails significant computational cost. For instance, fine-tuning a GPT-4 model for three epochs on a dataset with 50k rows and a five-column table costs approximately \$300 (OpenAI) or requires equivalent high-end GPU resources. Therefore, while the same fine-tuning schema can be applied to models like GPT-4 or LLaMA, it is not a cost-effective solution compared to the solutions considered here.
In summary, newer generation LLMs do not obviate the need for PAFT . Our method directly addresses the impedance mismatch between autoregressive LLMs and synthetic table generation—particularly by preserving functional dependencies through permutation-aware fine-tuning—a challenge that persists even with advanced models.
6. Conclusion
This work has brought LLMs closer to the goal of generating realistic synthetic datasets. By learning FDs and leveraging this information in the fine-tuning process, we are able to align the auto-regressive nature of LLMs with the ordering of columns necessary for generating quality synthetic data. While PAFT is quite broadly applicable by itself, it can be extended in several directions. First, what are other, perhaps more expressive, types of tabular constraints that can be utilized in the fine-tuning process? Second, what is the internal basis for regulating orders inside a transformer architecture and can we more directly harness it? Third, can we theoretically prove the (im)possibility of generating specific synthetic datasets by LLM architectures? Fourth, despite the numerous advantages of LLM in learning and generating tabular data, scalability remains an acknowledged challenge (Dinh et al., 2022; Borisov et al., 2022; Fang et al., 2024), encompassing concerns such as context window and training speed. And finally, privacy-preserving methods have been implemented in table generators based on GANs but remains understudied in LLM fine-tuning. These questions will be the focus of our future work.
Limitations. The row-wise generation cost of our method, particularly when employing fine-tuning, is affected by the dataset sample size and computational resources (GPU). Moreover, the capacity of PAFT to generate columns is affected by context window sizes. These limitations can be overcome by the newer generation of LLMs or by exploring partial row generation, i.e., generating a row in multiple steps using an LLM.